# Bootstrapping and autonomous tuning system for Si/SiGe multi-quantum dot devices

### v3

@@ IV BATIS: Experimental Demonstration
Following the gate-biased illumination, the fourth iteration of the global turn-on measurement was successful. The summary of all four runs is presented in Table 1. The resulting configuration of the QD device is shown in Fig. 5(a), where the pinch-off voltages on all the screening gates, excluding S2, and all of the reservoir gates are colored based on the learned gate voltage.
Accumulated Leakage Test. With the global accumulation confirmed, BATIS proceeds to the accumulated leakage test. A second leakage matrix is measured, following the same procedure as during the first leakage test. Like previously, BATIS finds that all resistances exceed 25 $\mathrm{M\SIUnitSymbolOhm}$. Since none of the connections leaked, only $35$ measurements following the matrix diagonal were taken.
@@
Screening and Reservoir Characterization The next step BATIS follows involves characterization of the screening and reservoir gates, which is important for dividing the induced 2DEG into three current channels. The results of these characterizations for all tested gates are presented in Fig. 6(b). The S2 gate borders all of the current channels and thus is not included here as it needs to be measured simultaneously on channels I1, I2, and I3. The data for the S2 pinch-off can be found in Fig. 6(c).
@@
Figure 6: Successful BATIS tune-up on a three-channel, six-QD device. (a) A schematic circuit diagram of the QD device shown in Fig. 1(a) depicting how BATIS understands the device. Numbered current paths are shown, as are the source voltage and current amplifier connections to ohmic contacts. This information is entered into the configuration file. (b) Pinch-off data for all reservoir and external screening gates grouped by measurement channel. The acquired signal for gate S1 extends over twice the voltage range of all other measurements. However, as the low-voltage data far from the transition region exhibits no strong features, the plotted voltage range is truncated for visualization purposes. (c) The results of the channel accumulation measurements for each current channel. The $\mathcal{F}$ axis represents the normalized voltage range applied to all of the finger gates, and the $\mathcal{S}$ axis is the normalized voltage range applied to the screening gates bordering a given channel. The finger gates for channel I1 in panel (a) are swept from 104-104 $\mathrm{mV}$ to $724$ $\mathrm{mV}$, the finger gates for channel I2 in panel (a) are swept from 91-91 $\mathrm{mV}$ to $727$ $\mathrm{mV}$, and the finger gates for channel I3 are swept from 83-83 $\mathrm{mV}$ to $729$ $\mathrm{mV}$. The screening gates are swept from $127$ $\mathrm{mV}$ to $427$ $\mathrm{mV}$, and 18-18 $\mathrm{mV}$ to $282$ $\mathrm{mV}$, and from $50$ $\mathrm{mV}$ to $350$ $\mathrm{mV}$ for channels I1, I2, I3, respectively. The S2 pinch-off curves are included for each channel. (d) Pinch-off data for all finger gates. BATIS has adjusted the trapped charge under the gates through biased illumination so that all pinch-off voltages occur in a reasonable range, in this case, between 200-200 $\mathrm{mV}$ and $500$ $\mathrm{mV}$. The color bar on the right corresponds to all plots in (b), (c), and (d).
@@
The voltage values are quite consistent between the sets of reservoir and screening gates intended to form the desired current channels, a feature that arises in part from the gate-biased illumination described above [34]. We note that for gate S1, BATIS acquired the signal over a $600$ $\mathrm{mV}$ range, i.e., twice as long compared to all other screening gates. This is due to poor fit scoring using the reduced voltage scan range, which caused BATIS to perform additional measurements with an increased sweep range to ensure sufficient confidence in the derived pinch-off value. Since the long tail of the signal has no features, the signal shown in Fig. 6(b) is truncated to be consistent with the remaining measurements for visualization purposes.
@@
The summary of the pinch-off voltages and the operating points selected by BATIS is presented in Table 2. A device in this configuration prepared for the channel accumulation stage is shown in Fig. 5(b).
@@
Table 2: The summary of the derived voltages for screening and reservoir gates as determined during their characterization stage by BATIS. The first column indicates the gate. The pinch-off voltages, VP, the guaranteed isolation voltages for screening gates, VI, and the operating points for reservoirs, VO, are presented in the next three columns. The gates are organized top-down in the order they were measured.
Gate VP ($\mathrm{mV}$) VI ($\mathrm{mV}$) VO ($\mathrm{mV}$)
@@
Current Channel Accumulation With all screening and reservoir gates characterized and candidate operating voltages determined, BATIS proceeds to form the 1D current channels. As discussed above in Sec. III.5, the algorithm for 1D current channel formation implemented in BATIS does not follow the typical, triangle-plot-based strategy. Rather, it sweeps each of the three outer screening gates, i.e., S1, S3, and S4, against all finger gates relevant to the given 1D current channel. The traditional triangle plots are along a different set of axes (e.g., S2 and S1, instead of all finger gates and S1) than those shown in Fig. 6(c). The advantage of measurements implemented in BATIS is faster convergence to the correct operating point for three well-formed and independent current channels.
[…]
If BATIS was unable to select an operating point after such a measurement, the next action would be to adjust all of the gates by using a 400 $\mathrm{mV}$ illumination. This adjustment can be found in the configuration file. After such an illumination, the entire device response can change because of crosstalk and the interdependence of the current paths. Thus, BATIS resumes autotuning from the global accumulation step.
The quad-QD device requires three current channels to be formed, as indicated with the white arrows in Fig. 6(a). In this demonstration, BATIS identified three different operating points, one for each 1D current channel, based on the data shown in Fig. 6(c). The summary of the screening and finger gates configuration for each operating point can be found in Table 3, and a collapsed visualization of these gate voltages is shown in Fig. 5(c).
[…]
@@
The accumulation plots in Fig. 6(c) reveal a dip in the current exiting $I_{3}$ when the current starts to flow in $I_{2}$, shown by the dark green shape imposed on top of the signal in $I_{3}$. This is due to a coupling between those two channels resulting from a shared ohmic connection to source the currents. Such a dip in the charge sensor readout is often called a shadow. Shadows increase the difficulty of detecting and forming isolated current channels, yet BATIS can handle this situation. In contrast, the accumulation plot for I1 is much less correlated with the current flow in the two charge sensor channels, but there is still some crosstalk arising from gate-to-2DEG cross capacitances.
Finger Gate Characterization The last stage of BATIS involves the determination of pinch-off voltages for the finger gates. One of the problems with analyzing pinch-off data is the presence of spurious Coulomb blockade oscillations, as discussed above and shown in the bottom row in Fig. 3(e). Such oscillations complicate fit-based analysis. Instead, BATIS uses thresholding and slope detection to determine the pinch-off voltage, searching in the increasing voltage direction (the reverse of how the signal was measured) to capture the first significant turn-on above the noise.
Table 4: The final pinch-off values, as determined by BATIS, for all the finger gates. The three main columns are associated with a set of finger gates belonging to a particular current channel on the device $\mathcal{F}[\mathrm{I}_{\ell}]$, with $\ell=1,2,3$. Each of the main columns consists of two sub-columns, with the left sub-column listing the finger gates belonging to the selected channel, and the right sub-column giving the pinch-off voltage VP determined by BATIS.
@@
Gate VP ($\mathrm{mV}$) Gate VP ($\mathrm{mV}$) Gate VP ($\mathrm{mV}$)
@@
This pinch-off voltage determination is performed individually for all of the finger gates of the QD device. The voltage range for each measurement is always the same and predefined in the configuration file. However, due to the Coulomb blockade oscillations, the automated search for the starting voltage for some gates is more challenging than for others, as is visible in the data shown in Fig. 6(d), and the added complexity of the curves informed the algorithm described above and implemented here. The resulting pinch-off values of all of these varying curves for all finger gates are presented in Table 4. This configuration is used in the final colorized SEM for the ending state of the algorithm shown in Fig. 5(d). This voltage configuration describes the proper operating point in high-dimensional voltage space. BATIS has taken the device from an unknown initial state to a set of properly defined 1D channels with confirmed finger gate operation and a full compilation of pinch-off voltages. At this point, autotuners designed for 1D devices can be used on each individual channel, if desired [9].
@@ V Summary and outlook
In this paper, we demonstrated an autonomous system for bootstrapping, automated testing, and characterization of gate-defined, accumulation-mode, undoped QD devices. This method, which we call BATIS, leverages knowledge from experiment and physical intuition—grounded in fundamental device physics—to electronically control navigation of the quantum device landscape. The prior knowledge about the QD device required to initialize BATIS is limited to the heterostructure type, the device geometry, the ohmic biases, and the device safety limits. The method proceeds in several steps, checking for device functionality as it moves forward. If irrecoverable device flaws are discovered, BATIS reports those issues and halts. Importantly, BATIS can recover from other non-idealities, such as undesirable amounts of trapped charge arising from fabrication or accident, by using illumination under gate-bias to modify in-situ the trapped charge and thus the device performance. This capability is important for resolving issues like charge disorder or incomplete turn-on, which may prevent functional behavior in otherwise operational devices.
@@
While much of the automated approach relies on existing experimental procedures, an innovation in BATIS is its method for 1D channel formation: rather than sweeping pairs of screening gates, BATIS performs a single 2D voltage scan between one screening gate and all intermediate finger gates. This strategy reduces the required measurement overhead to a single scan, resulting in substantial efficiency gains compared to conventional approaches.
@@
Importantly, the single-button-press testing and characterization process is demonstrated at $1.3$ $\mathrm{K}$, eliminating the need for a full $\mathrm{mK}$ cooldown for the device initialization. This result is particularly important as work advances towards “hot”, large-scale spin-based quantum processors. BATIS can also serve as a screening tool before cooling a dilution refrigerator to reach $\mathrm{mK}$ temperatures, if cold operation is desired.
@@
Previous studies have demonstrated device-agnostic algorithms for certain subsets of QD device-tuning tasks, such as setting QD topologies [47, 22] or charge-state classifications [14, 16]. These algorithms presuppose that device characterization and foundational tuning steps—such as determining gate functionality, defining the QD channel(s), or locating regions of interest for the coarse-tuning regime—have already been performed. BATIS helps to fill this gap, enabling bootstrapping of gate-defined devices in which current paths are entirely defined by gate voltages whose values cannot be known in advance.
@@ Acknowledgements.
We acknowledge Danielle Middlebrooks for helping with the automation. We acknowledge HRL Laboratories, LLC, for support and L.F. Edge for providing one of the Si/SixGe1-x heterostructures used in this work. This work was supported in part by ARO grant nos. W911NF-17-1-0274, W911NF-23-1-0110, and W911NF-24-2-0043. The authors gratefully acknowledge the use of facilities and instrumentation in the UW-Madison Wisconsin Center for Nanoscale Technology, which is partially supported by the Wisconsin Materials Research Science and Engineering Center (NSF DMR-2309000) and the University of Wisconsin-Madison. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright noted herein. Any mention of commercial products is for information only; it does not imply recommendation or endorsement by NIST.

### v4

@@ IV.3 Accumulated Leakage test
IV.4 Screening and reservoir characterization
@@
The next step in BATIS involves characterization of the screening and reservoir gates, which is important for dividing the induced 2DEG into three current channels. The results of these characterizations for all gates tested are presented in Fig. 6(b). Gate S2 borders all of the current channels and thus is not included here as it needs to be measured simultaneously on channels I1, I2, and I3. The data for the S2 pinch-off can be found in Fig. 6(c).
@@
Figure 6: Successful BATIS tune-up on a three-channel, six-QD device. (a) Schematic circuit diagram of the QD device shown in Fig. 1(a) depicting how BATIS understands the device. Numbered current paths are shown, as are the source voltage and current amplifier connections to Ohmic contacts. This information is entered into the configuration file. (b) Pinch-off data for all reservoir and external screening gates grouped by measurement channel. The acquired signal for gate S1 extends over twice the voltage range of all other measurements. However, as the low-voltage data far from the transition region exhibit no strong features, the plotted voltage range is truncated for visualization purposes. (c) Results of the channel accumulation measurements for each current channel. The $\mathcal{F}$ axis represents the normalized voltage range applied to all of the finger gates, and the $\mathcal{S}$ axis is the normalized voltage range applied to the screening gates bordering a given channel. The finger gates for channel I1 in (a) are swept from 104-104 $\mathrm{mV}$ to $724$ $\mathrm{mV}$, the finger gates for channel I2 in (a) are swept from 91-91 $\mathrm{mV}$ to $727$ $\mathrm{mV}$, and the finger gates for channel I3 are swept from 83-83 $\mathrm{mV}$ to $729$ $\mathrm{mV}$. The screening gates are swept from $127$ $\mathrm{mV}$ to $427$ $\mathrm{mV}$, from 18-18 $\mathrm{mV}$ to $282$ $\mathrm{mV}$, and from $50$ $\mathrm{mV}$ to $350$ $\mathrm{mV}$ for channels I1, I2, and I3, respectively. The S2 pinch-off curves are included for each channel. (d) Pinch-off data for all finger gates. BATIS has adjusted the trapped charge under the gates via biased illumination so that all pinch-off voltages fall within a reasonable range, in this case between 200-200 $\mathrm{mV}$ and $500$ $\mathrm{mV}$. The color bar on the right corresponds to all plots in (b)–(d).
@@
The voltage values are quite consistent between the sets of reservoir and screening gates intended to form the desired current channels, a feature that arises in part from the gate-biased illumination described above [34]. We note that for gate S1, BATIS acquired the signal over a $600$ $\mathrm{mV}$ range, i.e., twice as long as for all other screening gates. This is due to poor fit scores obtained with the reduced-voltage scan range, which prompted BATIS to perform additional measurements with an expanded sweep range to ensure sufficient confidence in the derived pinch-off value. Since the long tail of the signal has no features, the signal shown in Fig. 6(b) is truncated to be consistent with the remaining measurements for visualization purposes.
@@
A summary of the pinch-off voltages and the operating points selected by BATIS is presented in Table 2. A device in this configuration prepared for the channel accumulation stage is shown in Fig. 5(b).
@@
Table 2: Summary of the derived voltages for screening and reservoir gates determined during their characterization stage by BATIS. The pinch-off voltages, $V_{\mathcal{P}}$, the guaranteed isolation voltages $V_{\mathcal{I}}$ for screening gates, and the operating points $V_{\mathcal{O}}$ for reservoirs, are presented. The gates are organized top-down in the order they were measured.
Gate $V_{\mathcal{P}}$ ($\mathrm{mV}$) $V_{\mathcal{I}}$ ($\mathrm{mV}$) $V_{\mathcal{O}}$ ($\mathrm{mV}$)
@@
IV.5 Current channel accumulation
With all screening and reservoir gates characterized and candidate operating voltages determined, BATIS proceeds to form the 1D current channels. As discussed in Sec. III.5, the algorithm for 1D-current-channel formation implemented in BATIS does not follow the typical, triangle-plot-based strategy. Rather, it sweeps each of the three outer screening gates, i.e., S1, S3, and S4, against all finger gates relevant to the given 1D current channel. The traditional triangle plots are along a different set of axes (e.g., S2 and S1, instead of all finger gates and S1) than those shown in Fig. 6(c). The advantage of measurements implemented in BATIS is faster convergence to the correct operating point for three well-formed and independent current channels.
The automatic analysis of the regions and their respective boundaries is enabled by fitting a synthetic model to the experimental data, as explained in Ref. [46]. Additional criteria based on the local properties (e.g., estimated resistance and derivatives of the measurement) are used to optimize the selection of the final operating point $V_{\mathcal{O}}^{I_{\ell}}$ for $\ell=1,2,3$, among the numerous candidates in the operating region.
If BATIS was unable to select an operating point after such a measurement, the next action would be to adjust all gates using a 400 $\mathrm{mV}$ illumination. This adjustment can be found in the configuration file. After such an illumination, the entire device response can change because of crosstalk and the interdependence of the current paths. Thus, BATIS resumes autotuning from the global accumulation step.
The quadruple QD device requires three current channels to be formed, as indicated by the white arrows in Fig. 6(a). In this demonstration, BATIS identified three different operating points, one for each 1D current channel, on the basis of the data shown in Fig. 6(c). A summary of the configuration of the screening and finger gates for each operating point can be found in Table 3, and a collapsed visualization of these gate voltages is shown in Fig. 5(c).
The accumulation plots in Fig. 6(c) reveal a dip in the current exiting channel $I_{3}$ when the current starts to flow in channel $I_{2}$, shown by the dark green shape imposed on top of the signal in channel $I_{3}$. This is due to a coupling between those two channels resulting from a shared Ohmic connection to the current sources. Such a dip in the charge sensor readout is often called a “shadow.” Shadows increase the difficulty of detecting and forming isolated current channels, yet BATIS can handle this situation. In contrast, the accumulation plot for channel I1 shows much less correlation with the current flow in the two charge sensor channels, although some crosstalk still arises from gate-to-2DEG cross-capacitances.
IV.6 Finger gate characterization
The last stage of BATIS involves determining the pinch-off voltages for the finger gates. One of the problems with analyzing pinch-off data is the presence of spurious Coulomb blockade oscillations, as discussed above and shown in the third row in Fig. 3(e). Such oscillations complicate fit-based analysis. Instead, BATIS uses thresholding and slope detection to determine the pinch-off voltage, searching in the increasing voltage direction (the reverse of how the signal was measured) to capture the first significant turn-on above the noise.
Table 3: Screening and finger gate configuration summary for the three operating points identified by BATIS. Columns 2–4 provide the screening and finger gate operational configuration, $V^{I_{\ell}}_{\mathcal{O}}$ with $\ell=1,2,3$, for each of the operating points identified by BATIS to form channels I1, I2, and I3, respectively. The finger gates are grouped on the basis of the channel they belong to as $\mathcal{F}[\mathrm{I}_{1}]$, $\mathcal{F}[\mathrm{I}_{2}]$, and $\mathcal{F}[\mathrm{I}_{3}]$.
Gates $V^{I_{1}}_{\mathcal{O}}$ ($\mathrm{mV}$) $V^{I_{2}}_{\mathcal{O}}$ ($\mathrm{mV}$) $V^{I_{3}}_{\mathcal{O}}$ ($\mathrm{mV}$)
@@ IV.6 Finger gate characterization
Table 4: Final pinch-off values, as determined by BATIS, for all the finger gates. The three main columns are associated with a set of finger gates belonging to a particular current channel on the device $\mathcal{F}[\mathrm{I}_{\ell}]$, with $\ell=1,2,3$. Each of the main columns consists of two subcolumns, with the left subcolumn listing the finger gates belonging to the selected channel, and the right subcolumn giving the pinch-off voltage $V_{\mathcal{P}}$ determined by BATIS.
@@
Gate $V_{\mathcal{P}}$ ($\mathrm{mV}$) Gate $V_{\mathcal{P}}$ ($\mathrm{mV}$) Gate $V_{\mathcal{P}}$ ($\mathrm{mV}$)
@@
This pinch-off voltage determination is performed individually for each finger gate of the QD device. The voltage range for each measurement is always the same and predefined in the configuration file. However, because of the Coulomb blockade oscillations, the automated search for the starting voltage for some gates is more challenging than for others, as is visible in the data shown in Fig. 6(d), and the added complexity of the curves informed the algorithm described above and implemented here. The resulting pinch-off values of all these differing curves for all finger gates are presented in Table 4. This configuration is used in the final colorized scanning electron microscopy image for the ending state of the algorithm shown in Fig. 5(d). This voltage configuration describes the proper operating point in high-dimensional voltage space. BATIS has taken the device from an unknown initial state to a set of properly defined 1D channels, with confirmed fingergate operation and a full compilation of pinch-off voltages. At this point, autotuners designed for 1D devices can be used on each individual channel, if desired [9].
@@ V Summary and outlook
In this paper, we demonstrated an autonomous system for bootstrapping, automated testing, and characterization of gate-defined, accumulation-mode, undoped QD devices. This method, which we call BATIS, leverages knowledge from experiments and physical intuition—grounded in fundamental device physics—to electronically control the navigation of the quantum device landscape. By being rooted in heuristics and physical principles, the non-machine-learning implementation of BATIS provides a robust and interpretable baseline for developing and benchmarking future model-free strategies.
@@
The prior knowledge about the QD device required to initialize BATIS is limited to the heterostructure type, the device geometry, the Ohmic biases, and the device safety limits. The method proceeds in several steps, checking device functionality at each step. If irrecoverable device flaws are discovered, BATIS reports those issues and halts. Importantly, BATIS can recover from other nonidealities, such as undesirable amounts of trapped charge arising from fabrication or accidents, by using illumination under gate bias to modify in situ the trapped charge and thus the device performance. This capability is important for resolving issues such as charge disorder or incomplete turn-on, which may prevent functional operation in otherwise operational devices.
@@
While much of the automated approach relies on existing experimental procedures, an innovation in BATIS is its method for 1D channel formation: rather than sweeping pairs of screening gates, BATIS performs a single 2D voltage scan between one screening gate and all intermediate finger gates. This strategy reduces the required measurement overhead to a single scan, resulting in substantial efficiency gains compared with conventional approaches.
BATIS is built for scalability and extensibility. Its flexible architecture supports characterization and initialization of any accumulation-mode linear QD array and can be adapted to other QD platforms with minimal modification. For example, BATIS can replicate tuning routines developed for nanowires or GaAs/AlxxGa1x{1-x}As devices by simply skipping steps unnecessary for those architectures. In contrast, it is quite unlikely that any previously proposed algorithms could be easily adapted to the Si/SixGe1-x quadruple QD device considered in this work.
@@
Importantly, the single-button-press testing and characterization process is demonstrated at $1.3$ $\mathrm{K}$, eliminating the need for a full millikelvin cooldown for the device initialization. This result is particularly important as work advances toward “hot,” large-scale spin-based quantum processors. BATIS can also serve as a screening tool before cooling a dilution refrigerator to reach millikelvin temperatures, if cold operation is desired.
@@
Previous studies have demonstrated device-agnostic algorithms for certain subsets of QD-device-tuning tasks, such as setting QD topologies [47, 22] or charge-state classifications [14, 16]. These algorithms presuppose that device characterization and foundational tuning steps—such as determining gate functionality, defining the QD channel(s), or locating regions of interest for the coarse-tuning regime—have already been performed. BATIS helps to fill this gap, enabling bootstrapping of gate-defined devices in which current paths are entirely defined by gate voltages whose values cannot be known in advance.
@@ Acknowledgements.
We acknowledge Danielle Middlebrooks for helping with the automation. We acknowledge HRL Laboratories, LLC, for support and L.F. Edge for providing one of the Si/SixGe1-x heterostructures used in this work. This work was supported in part by ARO Grants No. W911NF-17-1-0274, No. W911NF-23-1-0110, and No. W911NF-24-2-0043. The authors gratefully acknowledge the use of facilities and instrumentation in the University of Wisconsin-Madison Wisconsin Center for Nanoscale Technology, which is partially supported by the Wisconsin Materials Research Science and Engineering Center (NSF Grant No. DMR-2309000) and the University of Wisconsin-Madison. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright noted herein. Any mention of commercial products is for information only; it does not imply recommendation or endorsement by NIST.
Data availability
The data that support the findings of this article are openly available [48].